# An approach to non-equilibrium statistical physics using variational Bayesian inference

### v1

@@ Acknowledgements
This paper more fully develops the theory outlined in a previous manuscript [1], and we are grateful to those listed there for their contributions to preliminary discussions on the topic. The account here is largely based on a series of talks given by the second named author at the Department of Philosophy of the University of Vienna and the Laufer Centre for Physical and Quantitative Biology, and the informal discussions that followed.
@@ 1 Introduction
In statistics, generative models are joint probability distributions that model the relationships between different variables—most often, observations and their causes, or data and a set of labels. In the study of open and interacting systems, the use of statistical models—capable of describing such relationships—seems to be particularly apt. One approach to making such models, using variational (approximate Bayesian) inference [2, 3], describes how a physical object reflects the properties of its environment and vice versa. Bayesian mechanics under the free energy principle (FEP) [4, 5] aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference), resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled.
@@
In this paper, we argue that generative models provide an especially useful foundation in scientific modelling generally, and especially in the study of systems exhibiting adaptation, morphogenesis, and other non-equilibrium phenomena. More specifically, we argue that generative modelling—premised on the FEP—is particularly apt for the practice of modelling the activity of modelling itself; for instance, in modelling sophisticated self-organising systems, where we write down a representation of the model that we assume is in play in prediction and action. We use the FEP to illustrate how implementing models which (i) formalise how one object keeps track of another, and (ii) codify these relationships, can increase the fidelity of our mathematical modelling. We identify a particular issue, which the FEP is able to uniquely address: the FEP keeps track of how a model depends on aspects of the thing that is modelled in a particularly insightful way.
Interestingly, because it incorporates this information about things that are coupled and their coupling, it has been suggested that the FEP conflates the metaphorical map (our mathematical model of the thing modelled) and the territory (the thing modelled). The so-called model reification or map-territory fallacy is a general critique of the practice of scientific modelling—in physics, principally due to Cartwright [6, 7]. Some have referenced this fallacy in relation to the FEP [8, 9]. In particular, it has been suggested that FEP-theoretic modelling conflates the metaphorical “map”—(i.e., the scientific model that scientists use to make sense of some real-world phenomenon)—and “territory,” i.e., the actual physical system that is being modelled. (See [10, 11, 12, 3] for related critical discussions of such claims.)
@@
According to this line of thinking, using Bayesian mechanics to claim that we can model objects as themselves engaging in inference about the statistics of their environment constitutes a case of model reification. The allegation is that, in describing the dynamics of physical objects as implementing a form of inference—as opposed to considering the inferential aspects of the explanation as pertaining to our scientific models of those objects—the FEP theorist mistakenly conflates their metaphorical “map” of the territory (i.e., the scientific FEP-theoretic or Bayesian mechanical model) with the “territory” (the real-world system itself). This constitutes a reification of our model by assuming some aspect of the model or “map” is a real feature of the physical world or “territory” [10, 11]. By clarifying that this potential difficulty does not directly and necessarily apply to FEP-theoretic modelling, we are able to turn the story on its head: namely, we point out that the FEP provides a set of ultimate explanatory constraints on what counts as a model (or map) of anything whatsoever that exists in the physical world.
@@
Here, we discuss the use of the FEP in statistics and statistical physics, and what advantages it offers; we then argue that there is no conflation of map and territory, i.e., of modelled system and scientific model, in the FEP—in fact, we describe how they are carefully distinguished. We argue that modelling the dynamics of an object such that it looks as if it models the system it is coupled to is not an instance of the map-territory fallacy. This argument gives way to a more general theme of the FEP, which is that capturing interactions and dependence relations between things using generative models is insightful. We discuss how using the FEP gives us a flexible model of the properties of the object that look like they perform inference, but crucially, without inherently reifying the inferential aspects of the model, attributing them to the real-world object. That is to say, a model of an object can always be written as modelling or inferring the structure of its environment. To make this argument, we appeal to two distinctions: one between the generative model (which is our scientific model) and the physical object being modelled (the real ‘thing’ in the world) [3]; and the other between the generative model and the variational density, which are often conflated in the secondary literature [13, 14].
@@
To sketch the argument: the FEP provides us with tools from dynamical systems and information theory, allowing us to model any coupled random dynamical systems as statistical estimators, i.e., as engaging in the statistical estimation of some quantity pertaining to the set of objects with which it interacts. Under the FEP, one object that is coupled to any other can thus be read as a representation of the coupled or estimated object, albeit of a deflationary, qualified sort (discussed in [15]; see also [13, 16, 3]). More formally, the FEP applies to generative models that contain a partition of the states or paths of a joint system (a so called “particular” partition, i.e., into objects or particles). This partition induces a statistical boundary called a Markov blanket [17], which captures conditional independence relations amongst subsets of the system [2, 18]. The FEP says that, when such Markov blanketed subsets of the system exist, those subsystems track [19] each others’ statistical properties (dynamically) across the blanket. Essentially, the FEP is a modelling framework which captures the following fact: models of coupled particles or factorised systems can be written as parameter estimators, and estimating the parameters of a distribution is (a form of) Bayesian inference. We can think of the FEP, metaphorically, as a map of the part of the territory that behaves as if it were a map—or looks as such to an external observer.
Accounts such as [10] have encouraged practitioners of the FEP to be explicit about this point. Viewing the theory this way, not only is there no conflation between our FEP-theoretic model of the system and the system itself, but in fact we arrive at a novel modelling method for coupled random dynamical systems. In particular, we do not claim that the objects in the system themselves must literally be performing inference. Rather, we describe the dynamics as doing inference, in the sense of being a statistical estimator whose states parameterise a variational density.
@@
We first present a conceptual, and then technical, introduction to the FEP and Bayesian mechanics—and the way it leverages the relationships encoded by a generative model to build a scientific model of the dynamics of physical objects using free energy gradients. This is possible by modelling the property of physical systems that they look as if they engage in modelling to an external observer. We then argue that there is no ambiguity between model and modelled real-world system in this formalism. We will examine possible reasons for mistakenly invoking the map-territory fallacy in this context, due mainly to the historically botched distinctions between the generative model and the variational density, on the one hand, and between the generative model and the so-called generative process, on the other.
We propose that the FEP and Bayesian mechanics provide a map of any possible map whatsoever—of, or held by—a physical system. (This includes the use of the FEP itself by a modeller.) We suggest that this leads, in a Wittgensteinian manner, to constraints on all possible maps, or models of physical things. This is meant in terms of logically necessary preconditions for sense-making, and thus the parsing of sensory streams on which cognition is contingent [15]. In particular, these constraints arise from what it means to be a map or model of a physical process at all.
@@
We remark that we will not claim to invalidate the map-territory fallacy itself. That is, we do not suggest that the map and the territory are never conflated; they obviously are in cases of genuine model reification (as discussed cogently by [11]). Models built using the FEP are still scientific models, and the properties of FEP theoretic model also run the risk of reification. Instead, we will address the claim that the FEP is limited in its effectiveness or engages in model reification because real physical things cannot be modelled as themselves modelling their environments by minimising free energy.
2 The free energy principle
2.1 Conceptually
In this subsection we will provide an up-to-date overview of the concepts behind the FEP. One can also see [3] for a discussion of the main ideas. See the following subsection for more mathematical details.
Conceptually, the FEP starts from a partitions an overall “system” into three component sets of states: the internal states $\mu$ of a ‘thing’ or ‘particle’, an environment (which possesses external states $\eta$, relative to the thing in question), and an interface separating but coupling the two (the so-called Markov blanket, $b$), which mediates the influence of $\eta$ on $\mu$ and vice-versa. The states of the system will be tuples $x\coloneqq(\eta,b,\mu)$, where $\eta$ and $\mu$ are states of two subsystems and $b$ is some locus of states such that $p(\eta\mid\mu,b)=p(\eta\mid b)$ and vice versa. “Particular states” are the internal states of each thing and their respective Markov blanket. Of note, the particular partition is akin to choosing a set of coarse-grained variables at a superordinate scale, associating each state or path at the subordinate scale to a particle at the superordinate scale (as either an internal or a blanket state), in terms of the relations of conditional independence that obtain between states or paths [20, 2].
More precisely, we will consider systems of interacting objects, such as an open random dynamical system coupled to its environment. Fixing a final time $\tau$, the joint system will be denoted by the time-indexed family of random variables $\{X\}_{t:\tau}$ taking values $x\in\mathcal{X}$ at each $t$, with a specific realisation of the system (meaning, a sample path or trajectory beginning at some fixed $X_{0}=x_{0}$) being the sequence $\gamma(x_{0})=\{x_{0},x_{1},\ldots,x_{\tau}\}$. We will define $\gamma(x_{0})\mapsto\gamma_{t}(x_{0})$ as the projection to the state at time $t$. Then $B_{t}$ is what we have called the Markov blanket in $\{X\}_{t:\tau}$ at time $t$, separating it into statistically distinct subsystems (which one can then argue do not mix into each other [18]). We will repeat this notation for
$\gamma_{t}(\eta_{0})=N_{t}=\eta\in\mathcal{N}$
$\gamma_{t}(b_{0})=B_{t}=b\in\mathcal{B}$
$\gamma_{t}(\mu_{0})=M_{t}=\mu\in\mathcal{M}$
as well as for conditional trajectories such as
$\gamma_{t}(\eta_{0}\mid B_{t}=b)=\eta_{b}\in\mathcal{N}\crossproduct\{b\}.$
Although we will not go into the theory of random dynamical systems, when equipped with a time evolution map acting on the sample space, these are all the ingredients necessary to define a random dynamical system [21], and this is implicit in our discussion; for details see the account in [14]. When $\{M\}_{t:\tau}$ is an open system coupled to an environment with states $\eta$, a value $\mu$ is called an internal state, and the subsystem $\{M\}_{t:\tau}$ is variously called an object, a thing, or a particle in the literature. This is the convention we will stick to.
We can guarantee that the dynamics of a system can be described as a flow on surprisal, and thereby, cast as implementing a form of inference, whenever there exists a “particular partition” of that system. At the core of the FEP is a mathematical statement about the behaviour of such systems. For a comprehensive overview, see [3]. In particular, the FEP allows us relate the dynamics of states or paths that are internal to the object, $\mu$, to the dynamics of external states, conditioned on boundary or blanket states $p(\eta\mid b)$—that is, understanding what goes on inside a subsystem by understanding what it is coupled to, as ‘seen’ via blanket states [5, 22].
The logic of the FEP can be summarised as follows. If there exists a particular partition of a system, then there exists a coupling between internal and external states or paths, given blanket states or paths; then, since the internal dynamics of a given object are coupled to other (external) subsets of the system—in such a way that they can be read as parameters for probability densities over the external subset—the FEP describes these interacting systems using variational inference. This is a core aspect of the FEP that distinguishes it from a straightforward application of density dynamics to the internal states per se.
This is where the interpretation of dynamics as a form of inference comes into play. Indeed, one way to state the FEP is:
Claim 1.
If a random dynamical system (i) is subject to a specific kind of sparse coupling or boundary condition, such that it possesses a Markov blanket, and (ii) has a subsystem $\{M\}_{t:\tau}$ under that blanket with characteristic states (in the sense that it occupies a pullback attractor), then $\{M\}_{t:\tau}$ can always be described as if it minimises a variational free energy functional of densities over states or trajectories of $\{N\}_{t:\tau}$ [14, 22].
Another statement one can make is that Bayesian mechanics formalises the truism that, on average, an object will do whatever it is that it does on average, given the sort of object it is. Behind this tautological statement, we find a formal apparatus that enables us to model a variety of dynamical systems in a parsimonious manner. We hope this will allow us to take some first steps towards a mechanistic description of how self-organisation manifests in physical systems with such pullback attractors [3].
Bayesian mechanics explains why every physical thing looks as if it were performing inference over the causes of its perturbations, forming (Bayes) optimal beliefs about those causes [22]. Bayesian mechanics begins from the minimisation of self-information (also called surprisal) or free energy, and allows us to write down mechanical theories (i.e., laws of motion) yielding the dynamics of systems on an attractor [19]. Surprisal is simply the negative log-probability of an observed outcome. To say that an object minimises (or follows a path of least) surprisal is to say that an object evolves towards the outcomes which are most probable, given the sort of thing that it is (meaning, its characteristic states of being, or characteristic trajectories through its state space).
Interestingly, this means that surprisal is suited to act as a kind of ‘ontological potential’ for the object, in the sense that the surprisal is equivalent to a constraint on what states and outcomes are accessible to the object; the trajectories of a physical system are random curves through the gradient of surprisal over the state space, which satisfy an equation of motion derived from the surprisal itself [4, 23]. As has been asserted in recent literature [19], one reading of the FEP would hold that the crux of the ubiquity of the FEP is that everything which is coupled to another thing holds a model of that thing to which it is coupled in the following, trivial sense: in virtue of this coupling, one subsystem estimates the parameters of the probability density over states (or trajectories) of the other subsystems (and vice versa). Using this property, they can be written as performing inference in the sense of variational Bayesian inference.
This coupling is sometimes discussed in terms of the ‘probabilistic beliefs’ that one object holds about its environment. Given that the term ‘belief’ is quite theoretically loaded (especially in philosophy), we ought perhaps to say ‘estimator,’ instead. Once more, our contention is that anything which estimates the parameters of a density is engaged in a form of Bayesian inference, by mathematical definition; i.e., it is mathematically indisputable that systems that estimate statistics represent, or encode, or otherwise hold probabilistic beliefs—whether the practitioner likes the word ‘beliefs’ or not, and irrespective of what commitments one has with respect to notions such as encoding and representation (e.g., [24]).
Generative models and variational methods allow us to licence these claims [25, 26]. Generative models are joint probability densities over causes and consequences which, here, are used to capture the dependencies between the variables constituting $\{X\}_{t:\tau}$, meaning the way that changes in some variables cause changes in others. Writing down a generative model for some real object in the world (the metaphorical territory) is building a metaphorical map of that territory (a formalisation or codification of the relations between the things in that territory).
Variational techniques comprise a family of inference techniques which allow a modeller to approximate some (usually intractable) posterior probability density under a (usually specified) generative model. In variational inference, we write down a parameterised probability density function called a variational density, which is used as an approximation to the posterior probability density. The quantity called variational free energy scores the difference between the variational density and the posterior over external states—here, a conditional density, given a particular blanket state. Hence, when some notion of the difference between them is zero, variational free energy can be used as a proxy for surprise, and we can approximate the posterior density; which is a form of parameterised Bayesian inference.
Note that any given object cannot change the generative model associated to that object. This follows because the object can only change the variational density encoded by its internal states. However, to explain these changes—under a variational principle of least action—one needs to specify the generative model under which the posterior is approximated. This means the generative model explains the dynamics of something that is observed in a particular way. This is usually expressed by saying that the generative model is entailed by the thing in question. That an object entails a generative model is consistent with statements like “an object is a model of its environment”: e.g., in the sense of the good regulator theorem [19]). This follows because the generative model just is a probabilistic description of the joint states an object and its environment occupy.
2.2 Mathematically
With that conceptual overview being stated, we will now provide some further mathematical details. Similar (contemporary) reviews can be found in [4, 5, 19, 27]. We will argue that coupled random dynamical systems partitioned under a Markov blanket naturally have the form of variational Bayesian inference problems, in the sense that we obtain the values of parameters from the minimisation of free energy or surprisal. This is what licences the claim that minimising free energy implements a form of inference, in which case the object models its environment. Ultimately this simply means that, in virtue of being coupled via blanket states to its environment, the object reflects the structure of, and data about (i.e., models), its environment—or, what it ‘believes’ that its environment is like (in the sense of being a statistical estimator, as defined above).
Appealing to the Laplace approximation (that if a density has a unique maximum, in a neighbourhood of the most likely state, then the density is approximately Gaussian with that mean), we will suppose all random variables and all trajectories are jointly Gaussian. We will call the conditionally expected state—that is, the expected state given some fixed blanket state—the conditional mode $\hat{\mu}_{b}$ or $\hat{\eta}_{b}$. Let $\mathcal{L}(\mu,b)=-\log p(\mu,b)$ be the surprisal of internal and blanket states, such that $\hat{\mu}_{b}=\operatorname*{\mathrm{arg\,min}}_{\mu}\mathcal{L}(\mu,b)$ and thus satisfies $\nabla_{\mu}\mathcal{L}(\mu,b)=0$. By the Freidlin–Wentzell theorem, we can argue the same is true of paths (see [19]). In that setting, the interpretation of surprisal as the fluctuations of a random dynamical system away from its expected path plays a crucial role in stochastic thermodynamics [28, 29].
The FEP states that, given the existence of a boundary, we can construct a map relating (the dynamics of) internal states to (those of) the environment across the boundary. In particular, if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma:\mathcal{M}\to\mathcal{N}$ such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, sending conditional modes to conditional modes [22, Lemma 4.3]. We construct this map by composing the natural projections to boundary states, $\bm{\mu}:\mathcal{B}\to\mathcal{M}$ and $\bm{\eta}:\mathcal{B}\to\mathcal{N}$, into
$\bm{\eta}\circ\bm{\mu}^{-1}:\mathcal{M}\to\mathcal{N}.$
We have assumed $\bm{\mu}$ is at minimum an injection, hence bijective on its image, from which we can take a left-cancellative pseudo-inverse. Under hypotheses on these maps, for example, that two blanket states which map to the same value of $\mu$ also map to the same value of $\eta$, this is a useful function, in the sense that it associates a unique pair of conditional modes to each blanket state. [^1] One can see [4] for a proof looking at the properties of solutions to certain linear systems of equations. Alternatively, note that this is evident from taking a tensor-Hom adjunct of the composition, $b\mapsto\xi(b,-)$, and asking that this map be an injection in the sense that when $i=j$, the map takes $\hat{\mu}_{i}=\hat{\mu}_{j}$ to equal images $\hat{\eta}_{i}=\hat{\eta}_{j}$ [22, pp 25]. Importantly, in the definition of $\sigma$, we can pass an arbitrary internal state in; its image will simply not be the conditional mode of the environment, and hence it will not parameterise a density consistent with $p(\eta\mid b)$.
We then postulate the existence of a family of probability densities parameterised by values of some known quantity, called the variational density $q(\eta;\hat{\eta})$. Generically this is constructed by taking a family of Gaussian densities with different modes. By the previous discussion, we can interpret this as the likely configuration of the environment, knowing the environment must be amenable to a certain internal state on average (given a corresponding blanket state), $q(\eta;\sigma(\mu_{b}))$. It is immediate that the probability density parameterised by the conditional mode $\hat{\eta}_{b}$ ought to be the same as the conditional probability $p(\eta\mid b)$. Moreover, if we assume the system is coupled by a particular partition, we naturally have $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$. Together, we have supposed that the object exists as the preimage of $\hat{\eta}_{b}$, or tends to do so, and hence, that the environmental states that are likely under a given conditional mode of the object ought to be equivalent to the probabilities over external states conditioned on that blanket state.
Because we are identifying a parameter value which both (i) has the greatest log-likelihood of being observed and (ii) fits a parameterised distribution over unobserved states to a distribution over unobserved states given observed states, this set up naturally has the structure of a variational inference problem [30]. As discussed, the quantity known as variational free energy scores the difference between the variational density and the true conditional density in terms of the relative entropy from one to the other, known as the KL divergence, added to the surprisal of internal and blanket states:

### v2

# A framework for the use of generative modelling in non-equilibrium statistical mechanics

@@ Acknowledgements
This paper more fully develops the theory outlined in a previous manuscript [1], and we are grateful to those listed there for their contributions to preliminary discussions on the topic. The account here is largely based on a series of talks given by the second named author at the Department of Philosophy of the University of Vienna, the Laufer Center for Physical and Quantitative Biology, the 2024 Conference on Complex Systems, and the informal discussions that followed. We thank the organisers and audiences of both. DARS acknowledges support from the Einstein Chair programme at the Graduate Centre of the City University of New York.
@@ 1 Introduction
In statistics, generative models are joint probability distributions that model the relationships between different variables—most often, observations and their causes, or data and a set of labels. In the study of open and interacting systems, the use of statistical models—capable of describing such relationships—seems to be particularly apt. One approach to making such models, using variational (approximate Bayesian) inference [2, 3, 4], describes how a physical object reflects the properties of its environment and vice versa. In this paper, we argue that generative models provide an especially useful foundation in scientific modelling generally, and especially in the study of systems exhibiting adaptation, morphogenesis, and other non-equilibrium phenomena. More specifically, we argue that generative modelling—premised on the use of variational inference to model the causes of the signals that a coupled system is receiving [5]—is particularly apt for the practice of modelling the activity of modelling itself; for instance, in modelling sophisticated self-organising systems, where we write down a representation of the model that we assume is in play in prediction and action. We use the ensuing variational free energy principle to illustrate how implementing models which (i) formalise how one object keeps track of another, and (ii) codify these relationships, can increase the fidelity of our mathematical modelling.
Our proposed free energy principle (FEP) has been explored in some depth in prior literature, from the viewpoint of modelling control systems using the property that they generally look as if they track or infer the statistics of perturbations to their setpoints. In this paper we will make a more general case for its utility in statistical mechanics at large. After reviewing prior work on the FEP, we identify a particular issue that the FEP is able to uniquely address: the FEP keeps track of how a model depends on aspects of the thing that is modelled in a particularly insightful way, offering advantages in terms of tractability and explanation, by using gradients of variational free energy to describe the dynamics of interacting systems.
@@
Interestingly, because it incorporates this information about things that are coupled and their coupling, it has been suggested that our FEP conflates the metaphorical map (our mathematical model of the thing modelled) and the territory (the thing modelled). The so-called model reification or map-territory fallacy is a general critique of the practice of scientific modelling—in physics, principally due to Cartwright [6, 7]. Some have referenced this fallacy in relation to prior work on the FEP [8, 9]. In particular, it has been suggested that FEP-theoretic modelling conflates the metaphorical “map”—(i.e., the scientific model that scientists use to make sense of some real-world phenomenon)—and “territory,” i.e., the actual physical system that is being modelled. (See [10, 11, 12, 3] for related critical discussions of such claims.) According to this line of thinking, using the FEP to claim that we can model objects as themselves engaging in inference about the statistics of their environment constitutes a case of model reification. The allegation is that, in describing the dynamics of physical objects as implementing a form of inference—as opposed to considering the inferential aspects of the explanation as pertaining to our scientific models of those objects—the FEP theorist mistakenly conflates their metaphorical “map” of the territory (i.e., the scientific, FEP-theoretic model) with the “territory” (the real-world system itself). This constitutes a reification of our model by assuming some aspect of the model or “map” is a real feature of the physical world or “territory” [10, 11]. By clarifying that this potential difficulty does not directly and necessarily apply to FEP-theoretic modelling, we are able to turn the story on its head: namely, we point out that the FEP provides a set of ultimate explanatory constraints on what counts as a model (or map) of any thing that exists in the physical world.
@@
After describing the advantages of this FEP we will then argue that there is no conflation of map and territory, i.e., of modelled system and scientific model, in the FEP—in fact, we describe how they are carefully distinguished. We argue that modelling the dynamics of an object such that it looks as if it models the system it is coupled to is not an instance of the map-territory fallacy. This argument gives way to a more general theme of the FEP, which is that capturing interactions and dependence relations between things using generative models is insightful. We discuss how using the FEP gives us a flexible model of the properties of the object that look like they perform inference, but crucially, without inherently reifying the inferential aspects of the model, attributing them to the real-world object. That is to say, a model of an object can always be written as if it were modelling or inferring the structure of its environment. To make this argument, we appeal to two distinctions: one between the generative model (which is our scientific model) and the physical object being modelled (the real ‘thing’ in the world) [3]; and the other between the generative model and the variational density, which are often conflated in the secondary literature [13, 14].
@@
To sketch the argument: the FEP provides us with tools from dynamical systems and information theory, allowing us to model any coupled random dynamical systems as statistical estimators, i.e., as engaging in the statistical estimation of some quantity pertaining to the set of objects with which it interacts. Under the FEP, one object that is coupled to any other can thus be read as a representation of the coupled or estimated object, albeit of a deflationary, qualified sort (discussed in [15]; see also [13, 16, 3]). More formally, the FEP applies to generative models that contain a partition of the states or paths of a joint system (a so called “particular” partition, i.e., into objects or particles). This partition induces a statistical boundary called a Markov blanket [17], which captures conditional independence relations amongst subsets of the system [2, 18]. The FEP says that, when such Markov blanketed subsets of the system exist, those subsystems track [19] each others’ statistical properties (dynamically) across the blanket. Essentially, the FEP is a modelling framework which captures the following fact: models of coupled particles or factorised systems can be written as parameter estimators, and estimating the parameters of a distribution is a form of Bayesian inference. We can think of the FEP, metaphorically, as a map of the part of the territory that behaves as if it were a map—or looks as such to an external observer. Accounts such as [10] have encouraged practitioners of the FEP to be explicit about this point. Viewing the theory this way, not only is there no conflation between our FEP-theoretic model of the system and the system itself, but in fact we arrive at a novel modelling method for coupled random dynamical systems. In particular, we do not claim that the objects in the system themselves must literally be performing inference. Rather, we describe the dynamics as doing inference, in the sense of being a statistical estimator whose states parameterise a variational density.
@@
The structure of the paper is as follows. We first present a conceptual introduction to the FEP—and the way it leverages the relationships encoded by a generative model to build a scientific model of the dynamics of physical objects using free energy gradients. This is possible by modelling the property of physical systems that they look as if they engage in modelling to an external observer. We then argue that there is no ambiguity between model and modelled real-world system in this formalism. We will examine possible reasons for mistakenly invoking the map-territory fallacy in this context, due mainly to the historically botched distinctions between the generative model and the variational density, on the one hand, and between the generative model and the so-called generative process, on the other. In summary we propose that the FEP provides a map of any possible map whatsoever—of, or held by—a physical system. (This includes the use of the FEP itself by a modeller.) We suggest that this leads to constraints on all possible maps, or models of physical things, arising from what it means to be a map or model of a physical process at all. In the fashion of Jaynes, who proposed that doing statistical mechanics is doing inference about the probabilities of microstates given knowledge of macrostates [20, 21], the FEP is a natural expression of statistical physics as a process of making models.
@@
We remark that we will not claim to invalidate the map-territory fallacy itself. That is, we do not suggest that the map and the territory are never conflated; they obviously are in cases of genuine model reification (as discussed cogently by [11]). Models built using the FEP are still scientific models, and the properties of FEP-theoretic models also run the risk of reification. Instead, we will address the claim that the FEP is limited in its effectiveness or engages in model reification because real physical things cannot be modelled as themselves modelling their environments by minimising free energy.
2 Summary remarks on the free energy principle
Here we will describe the proposed free energy principle (also directing the reader to prior literature such as [19, 3, 4] for a more extended look). The FEP aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference) [22], resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled [3, 23, 19]. Formulated most generally, the FEP says that the average path taken by an object that is sparsely coupled to its environment is characterised by a Lyapunov function—a function that decreases monotonically along the trajectories of a (deterministic) dynamical system. The utility of Lyapunov functions is that they guarantee stability, allowing us to understand the dynamical system as flowing on the gradient of the Lyapunov function. In the FEP, this Lyapunov function is surprisal or variational free energy, depending on the setting. In other words, to understand the claim that the FEP models a system as doing inference, it is crucial to understanding the ‘inference’ component of surprisal minimisation. Surprisal is simply the negative log probability. Variational free energy is a tractable (i.e., easily computed) upper bound on surprisal, where minimising variational free energy allows us to infer a target density by approximating it with a variational density. The surprisal being considered is, much like probability, surprisal given a model, in that it quantifies how surprising it would be for an observer—equipped with a model—to find the system (such as itself) in some non-system-like state or configuration. Minimising surprisal implies that the variational free energy is minimised, (i.e., the object-as-model is a good estimator of its environment [19]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
We consider a particle (also referred to as a ‘system’) with states $\mu$ coupled to another system with states $\eta$ (henceforth, the ‘environment’) through a set of variables $B$ with values $b$, and a steady state distribution $p(\eta,b,\mu)$ such that $\mu$ is conditionally independent of $\eta$ given $B$. After Pearl [17] we refer to $B$ as the Markov blanket of the system. We then postulate a conditional density over external states, $q(\eta;\hat{\eta}_{b})$, where we have parameterised the variational density by the conditional mode $\hat{\eta}_{b}$ of the environment. For convenience, we assume a Laplace approximation to the posterior, meaning that in a neighbourhood of its peak, the posterior is approximately Gaussian, and hence uniquely determined by its mean and variance. This guarantees that finding the right value of $\hat{\eta}_{b}$ is sufficient for $q$ to equal the posterior almost surely. More generally, under the maximum entropy assumption of a Gibbs density, any constraint function suffices. The crucial idea is that the dynamics of a system does inference when written as minimising a free energy functional, which is itself the consequence of reading surprisal as a Lyapunov function for those dynamics, following e.g. Freidlin–Wentzell. In turn, the dynamics of a system minimise surprisal or free energy whenever that system is equipped with a Markov blanket; and hence, the dynamics of a system with a particular partition can always be written as a gradient flow on surprisal. This gradient flow is what we identify as actually being the process (i.e., the performance) of inference. Explicit examples of this can be found in [24, 25, 19].
More specifically, the FEP says that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary [23, Lemma 4.3]. In that case, we can map these parameters to $\sigma(\hat{\mu}_{b})$, meaning that for as long as the particle is at state $\hat{\mu}_{b}$, the particle models its environment by estimating the parameters of its distribution. Ultimately this simply means that, in virtue of being coupled (via blanket states) to its environment, the particle reflects data about its environment—or, what it believes that its environment is like. Under the assumption that a particle which exists in an environment models its environment, free energy is implicitly a measurement of coherence. If the particle stops modelling its environment in a particular way (i.e., for a particular $\sigma$), it must have ceased to exist with the mean that $\sigma$ was constructed for; conversely, if a particle ceases to exist, it will stop modelling its environment.
Since the dynamics towards the mode induce a gradient flow on a divergence between the estimated parametric distribution and the true distribution, the FEP moreover proposes that these dynamics (in virtue of the coupling) can be modelled as a variational inference process, following a gradient flow on the variational free energy